Change a speed of 72.4 miles per hour to its equvalent in metres per second. (Round your answer to one decimal place)
1
Expert's answer
2020-09-28T08:10:24-0400

Let's convert miles per hour to metres per second:


"v = (72.4 \ \dfrac{mi}{h}) \cdot (\dfrac{1609 \ m}{1 \ mi}) \cdot (\dfrac{1 \ h}{3600 \ s}) = 32.4 \ \dfrac{m}{s}."

Answer:

"v = 32.4 \ \dfrac{m}{s}."